# The Curious Case of Arbitrariness in Machine Learning

### v1

@@ 4.1. Multiplicity and Uncertainty
  • Multiplicity is better suited to explore alternative interpretations. As we’ve discussed, choosing among different learned interpretations can introduce arbitrariness. Multiplicity, particularly Rashomon sets, enables exploration of these alternative interpretations. Examples include personalization with model mixtures, combining multiple models, homogenization concerns, etc. (Ma et al., 2020; Creel and Hellman, 2022; Wu et al., 2024; Bommasani et al., 2022; Black et al., 2021; Laberge et al., 2023; Breiman, 2001; Long et al., 2024)
@@
Note that our recommendations paint a broad picture of when the lens of multiplicity or uncertainty could be useful, but these are intended only as guidelines and deviation from these may be warranted in specific contexts.
4.2. Multiplicity and Bias-Variance Decomposition
@@
Another common measure of arbitrariness in decisions is the ‘variance’ from the bias-variance decomposition. Error in machine learning is often categorized into three parts: irreducible error, bias, and variance (Domingos, 2000; Kong and Dietterich, 1995; BREIMAN, 1996). This is known as the bias-variance decomposition. The terms ‘bias’ and ‘variance’ describe how well the model approximates the underlying distribution, i.e., approximation complexity, where ‘bias’ captures how well the chosen model fits the given data, while ‘variance’ reflects the model’s sensitivity to variations in the dataset.
@@
Understandably, at first glance, sensitivity to variations in the dataset—captured by ‘variance’—might seem like a natural way to measure arbitrariness in decision-making. However, bias-variance decomposition is typically confined to the analysis of a single model, focusing only on sensitivity to variations in the underlying dataset. While this framework provides valuable insights into a model’s behaviour—guiding preferences for models with lower bias or variance—it does not address the broader arbitrariness introduced by the entire learning pipeline. Multiplicity, in contrast, enables comparisons across different models generated through the pipeline. Formally,
(8) $Error=IrreducibleError+Bias+{}$ $\mathbf{Var_{D}[f_{\theta}(x)]}$  (Domingos, 2000; Kong and Dietterich, 1995; BREIMAN, 1996)
(9) $PredictiveMultiplicity={}$ $\mathbf{Var_{\theta\sim{\textrm{Prob}}_{R}(\theta|D)}[f_{\theta}(x)]}$  (Heljakka et al., 2022; Long et al., 2024; Hamman et al., 2024)
@@
where $f_{\theta}(.)$ is the predictive function for the learned parameter value $\theta$, and $Var[.]$ represents variance. We use a metric of predictive multiplicity that also uses variance to quantify multiplicity, facilitating clearer comparisons (Heljakka et al., 2022; Long et al., 2024; Hamman et al., 2024). While the concept of ‘variance’ from bias-variance decomposition focuses on a single model’s sensitivity to variations in the dataset, multiplicity, in contrast, captures changing predictions across multiple models. As we will explore further in §5, multiplicity allows us to navigate through multiple candidate models and even identify those with desirable bias-variance properties–such as ‘simpler’ models, often associated with lower variance (Semenova et al., 2024; Liu et al., 2022; Rudin et al., 2024; Semenova et al., 2022; Coupkova and Boutin, 2024).
We showed that multiplicity is neither universally redundant nor superior when compared to uncertainty or bias-variance decomposition. Instead, we offer guidance on when each approach is most appropriate. It is also crucial to recognize the common approximations made in the literature and to explicitly acknowledge their assumptions. For instance, predictive multiplicity has been used in the literature to measure epidemic uncertainty (Ali et al., 2021; Shahbab and Zaman, 2024; Brunet et al., 2022). This stems from the intuition that epistemic uncertainty is often estimated using Bayesian neural networks, whose behaviour can resemble Rashomon sets. However, this approximation is only valid if the models are calibrated, meaning they already account for aleatoric uncertainty (Hüllermeier and Waegeman, 2021). Furthermore, it is important to distinguish between multiplicity within the Rashomon set and uncertainty (of any kind) across the entire set of models, as discussed earlier. Thus, using multiplicity as a proxy for epistemic uncertainty is an approximation which may fail in situations with high aleatoric uncertainty or when the distribution over the Rashomon set is significantly different from that on the entire set of models.
@@
The interplay between multiplicity, uncertainty, and bias-variance decomposition remains complex, underscoring the need for further research to better understand and effectively utilize these concepts.
5. Exploring Alternative Interpretations with Multiplicity
@@
Now that we’ve discussed the formalization of multiplicity, let’s delve into its real-world implications. One of the biggest advantages of multiplicity lies in its ability to uncover and explore various ‘good’ learned interpretations (Black et al., 2022; Rudin et al., 2024). When multiple interpretations exist, it is reasonable to expect that some of them may exhibit certain desirable properties, such as better fairness, robustness, interpretability, etc. Additionally, moving beyond a single-model paradigm also opens the door to aggregating insights from multiple models. In this section, we study how multiplicity facilitates such exploration and its broader implications. Our discussion builds on what Black et al. (2022) refer to as the aggregate-level effects of multiplicity, while incorporating insights from more recent developments in the field.
5.1. Searching Instead of Optimizing
@@
Machine learning often involves tackling complex optimizations, including two commonly seen hierarchical optimization problems: bi-level optimization and constrained optimization. Bi-level optimization refers to scenarios where one optimization problem depends on variables governed by another nested optimization (Sinha et al., 2017; Zhang et al., 2024). A classic example is hyperparameter optimization. Constrained optimization, on the other hand, involves solving the optimization problems under specific constraints (Bertsekas, 2014; Gallego-Posada, 2024; Goodfellow et al., 2014). Examples include enforcing constraints of fairness, robustness, safety, etc. during model training (Gallego-Posada, 2024; Dwork et al., 2012). These classes of optimization problems can be notoriously difficult to solve. Challenges arise from the complexity of formalizing certain constraints, the difficulty of creating continuous and differentiable relaxations, the absence of closed-form solutions, or evolving client requirements, among others.
@@
Interestingly, multiplicity offers a practical workaround: if you can’t optimize, search for it! Brute-force strategies of searching through various potential solutions are well-established in both bi-level optimization and constrained optimization literature (Sinha et al., 2017; Zhang et al., 2024; Gallego-Posada, 2024). Though more efficient reductions and mathematical alternatives are preferred when feasible, they are not always possible. In such cases, searching through potential solutions to find the best fit becomes a viable strategy, and multiplicity plays a pivotal role in this search (see Figure 3). Multiplicity has been used in the literature to find models with lower bias (Black et al., 2024b; Simson et al., 2024; Ganesh et al., 2024; Islam et al., 2021), smaller model complexity (Semenova et al., 2022, 2024; Boner et al., [n. d.]; Rudin et al., 2024; Doctor et al., 2024; Coupkova and Boutin, 2024), better explanations (Sun et al., [n. d.]; Veran et al., 2023), improved generalizability (Li et al., 2024), and the ability to allow personalization (Ma et al., 2020). Beyond simply enumerating the Rashomon set to search for better models, several recent works have also shown how the visualization of the Rashomon set can empower developers to select models that meet their specific requirements (He and Shaposhnik, 2023; Eerlings et al., 2024; Luyten et al., 2024).
@@
Figure 3. The role of multiplicity in brute-force search for bi-level and constrained optimization problems in machine learning.
5.2. Ensembles and More
@@
Selecting the “best” model may not always be recommended, particularly when no single interpretation of the data can be optimal. Instead, combining insights from multiple models can often yield better results in such scenarios. Techniques like prediction ensembling, or bagging, have long been a central recommendation for stability in machine learning (Breiman, 1996; Dietterich, 2000; Long et al., 2024). Many methods in the multiplicity literature have capitalized on combining various forms of information from models in the Rashomon set. While literature in this direction primarily focuses on aggregating model explanations to create more stable and reliable explanations (Marx et al., 2023; Black et al., 2021; Jiang et al., 2024a; Laberge et al., 2023; Ley et al., 2023; Sreedharan et al., 2018; Baniecki et al., 2024; Gunasekaran et al., 2024; Mitruț et al., 2024; Hibbeln et al., 2024; Donnelly et al., 2023; Fisher et al., 2019; Dong and Rudin, 2020; Kobylińska et al., 2024; Zuin et al., 2021; Hamamoto and Egi, 2021; Kowal, 2022; Li et al., [n. d.]; Smith et al., 2020), other works have also shown the benefits of aggregating fairness scores (Coston et al., 2021), individual probabilities (Roth et al., 2023), or regression analysis to discover causality (Ueki and Kawasaki, 2013). These techniques demonstrate the value of leveraging multiplicity not just to select a single best model but instead to combine multiple learned interpretations.
5.3. Hacking Metrics with Multiplicity
@@
While multiplicity can enable the discovery of better models, it also introduces risks, particularly the potential of exploiting these search methods to circumvent regulatory requirements and interventions. This is more prevalent when broad principles, such as fairness, are reduced to specific benchmarks or metrics (Black et al., 2024a). By leveraging multiplicity, it becomes possible to ‘hack’ these metrics, producing models that meet the specified criteria without truly adhering to the underlying principles the metrics are intended to enforce (Coker et al., 2021; Cooper et al., 2021; Møllersen and Holsbø, 2023). Several studies in the literature have shown that such a search can indeed produce models capable of “regulatory-washing”, being able to manipulate explanations (Sharma et al., 2024; Aïvodji et al., 2021; Shahin Shamsabadi et al., 2022) and fairness scores (Forde et al., 2021; Black et al., 2024a; Ganesh et al., 2024). Such manipulation can also occur unintentionally-a result of overfitting to a given metric (Black et al., 2024a; Cooper et al., 2021; Ganesh, 2024)-underscoring the need for vigilance against the misuse of multiplicity and calling for a more robust operationalization of regulatory frameworks (Black et al., 2024a).
6. Multiplicity and Responsible AI
@@
We now move to placing multiplicity in the broader landscape of responsible AI. In this section, we examine the two key concerns for individuals originating from multiplicity, i.e., arbitrariness in model selection and outcome homogenization. Again, our discussion here builds on what was termed as the individual-level effects of multiplicity by Black et al. (2022), while also focusing on topics that have gained more interest recently, such as homogenization.
6.1. Arbitrariness as a Responsible AI Concern
@@
Arbitrary decisions in an automated system can be deeply concerning, particularly in critical domains such as law and medicine, where they can have direct and lasting impacts on human lives (Black et al., 2022; Gomez et al., 2024; Sokol et al., 2024; Barocas et al., 2023; D’Amour et al., 2022; Watson-Daniels et al., 2024). Borrowing an analogy from Gomez et al. (2024), imagine a judge deciding legal cases by flipping a coin. While this may seem extreme, it demonstrates how machine learning models can have arbitrariness embedded in them due to an analogous coin flip done by the developer during model design. This aligns closely with our ICA framework (§2.1), where we discuss the arbitrary choices by the developers that can contribute to multiplicity.
@@
There are contexts where a degree of “controlled randomness” may be acceptable—or even necessary (§6.2). However, arbitrariness is a significant concern in scenarios where individuals lack access to other ‘equivalent opportunities’ (Barocas et al., 2023). For instance, in hiring, some level of arbitrariness may be acceptable, or even useful to deal with the concerns of homogenization (§6.2). This is because individuals looking for a job often seek multiple opportunities across different companies, increasing their chances of being hired elsewhere. In contrast, domains such as law or medicine typically involve singular, high-stakes decisions with no equivalent alternatives. In such critical situations, the presence of arbitrariness raises serious concerns for the responsible deployment of machine learning models.
@@
As explored earlier, the negative effects of multiplicity extend beyond the arbitrariness of just the final prediction or decision (§2.2). For instance, multiplicity in counterfactual explanations can impact the stability and validity of algorithmic recourse (Sreedharan et al., 2018; Hancox-Li, 2020; Hasan and Talbert, 2022; Leventi-Peetz and Weber, 2022; Baniecki et al., 2024; Hamman et al., 2023; Leofante et al., 2023; Kobylińska et al., 2024; Jiang et al., 2024a). The feasibility or nature of recourse might hinge on arbitrary design choices made during model development. These decisions can have real-world implications; for instance, a recourse option provided by one model may become invalid if the model is updated, invalidating previous efforts. This inconsistency raises both ethical and legal concerns (Kobylińska et al., 2024; De Toni et al., 2024).
@@
This arbitrariness becomes even more problematic when it disproportionately impacts different individuals, particularly harming underrepresented demographics. As hinted earlier, a significant source of arbitrariness is the model’s lack of ability to learn the underlying distribution. In modern machine learning, underrepresented groups often face these information gaps, which can manifest as data scarcity due to limited historical records or a lack of understanding of cultural context for how data relates to predictions (Barocas et al., 2023; Hooker, 2021; Buolamwini and Gebru, 2018; Birhane, 2022; Geirhos et al., 2020). Research has consistently shown that such disparities exacerbate existing inequalities—whether through arbitrariness, uncertainty, or multiplicity (Farnadi et al., 2024). The resulting disproportionate harms across groups highlight the pressing need to address arbitrariness in critical domains (Gomez et al., 2024; Cooper et al., 2024; Ali et al., 2021).
@@
Impact of Other Responsible AI Constraints on Multiplicity: We saw that multiplicity is a critical consideration in responsible AI. Interestingly, several works have also explored how multiplicity interacts with other pillars of responsible AI. For instance, studies have shown that imposing fairness constraints can inadvertently increase multiplicity (Long et al., 2024; Cavus and Biecek, 2024a). However, Long et al. (2024) argue that multiplicity stands outside the fairness-utility trade-off, meaning improvements in fairness do not have to entail increased multiplicity. They demonstrate that multiplicity can often be reduced through techniques like ensembling, while maintaining fairness. Similarly, Kulynych et al. (2023) explore the interaction between differential privacy and multiplicity, revealing that introducing privacy constraints tends to increase multiplicity. These findings underscore the complex interplay between multiplicity and other responsible AI principles, highlighting the need for further research to understand and navigate these trade-offs effectively.
6.2. Multiplicity and Homogenization

### v2

@@ 4.1 Multiplicity and Uncertainty
  • Multiplicity is better suited to explore alternative interpretations. Choosing among different learned interpretations can introduce arbitrariness. Multiplicity, particularly Rashomon sets, enables exploration of these alternative interpretations. Examples include personalization with model mixtures, combining multiple models, homogenization concerns, etc. (Ma et al. 2020; Creel and Hellman 2022; Wu, Black, and Chandrasekaran 2024; Bommasani et al. 2022; Black, Leino, and Fredrikson 2021; Laberge et al. 2023; Breiman 2001; Long et al. 2024)
@@
Note that our recommendations paint a broad picture of when the lens of multiplicity or uncertainty could be useful, but these are intended only as guidelines, and deviation from these may be warranted in specific contexts.
4.2 Multiplicity and Variance
@@
Another measure of arbitrariness in decisions is the ‘variance’ from the bias-variance decomposition. Error in machine learning is categorized into three parts: irreducible error, bias, and variance (Domingos 2000; Brieman 1996). The terms ‘bias’ and ‘variance’ describe how well the model approximates the underlying distribution, i.e., approximation complexity, where ‘bias’ captures how well the chosen model fits the given data, while ‘variance’ reflects the model’s sensitivity to variations in the dataset.
@@
At first glance, ‘variance’ might seem like a natural way to measure arbitrariness in decision-making. However, bias-variance decomposition is typically confined to a single model, focusing only on sensitivity to the underlying dataset. While this is valuable, guiding preferences for models with lower bias or variance, it does not address the broader arbitrariness introduced by the entire learning pipeline. Multiplicity, in contrast, enables comparisons across different models generated through the pipeline. Formally,
$Error=IrreducibleError+Bias+{}\mathbf{Var_{D}[f_{\theta}(x)]}$ (7)
$PredictiveMultiplicity={}\mathbf{Var_{\theta\sim{\textrm{Prob}}_{R}(\theta|D)}[f_{\theta}(x)]}$ (8)
@@
where $f_{\theta}(.)$ is the predictive function for the learned parameter value $\theta$, and $Var[.]$ represents variance. We use a metric of predictive multiplicity that also uses variance to quantify multiplicity, facilitating clearer comparisons (Heljakka et al. 2022; Long et al. 2024; Hamman et al. 2024). While the concept of ‘variance’ from bias-variance decomposition focuses on a single model’s sensitivity to variations in the dataset, multiplicity, in contrast, captures changing predictions across multiple models. Black, Raghavan, and Barocas (2022) provides a similar comparison, bounding the expected disagreement between two models (potentially quantifying multiplicity) using the expected variance.
@@
We showed that multiplicity is neither universally redundant nor superior when compared to uncertainty or bias-variance decomposition. Instead, we offer guidance on when each approach is most appropriate. The interplay between multiplicity, uncertainty, and bias-variance decomposition remains complex, underscoring the need for further research to better understand and effectively utilize these concepts.
5 Exploring Alternative Interpretations
@@
Now that we’ve discussed the formalization of multiplicity, let’s delve into its real-world implications. One of the biggest advantages of multiplicity lies in its ability to explore various ‘good’ learned interpretations (Black, Raghavan, and Barocas 2022; Rudin et al. 2024). When multiple interpretations exist, it is reasonable to expect that some of them may exhibit certain desirable properties, such as better fairness, robustness, interpretability, etc. In this section, we study how multiplicity facilitates such exploration and its broader implications. Our discussion builds on what Black, Raghavan, and Barocas (2022) refer to as the aggregate-level effects of multiplicity, while incorporating insights from more recent developments in the field.
5.1 Searching Instead of Optimizing
@@
Machine learning often involves tackling complex optimizations, including two common hierarchical optimization problems: bi-level optimization and constrained optimization. Bi-level optimization refers to scenarios where one optimization problem depends on variables governed by another nested optimization (Sinha, Malo, and Deb 2017; Zhang et al. 2024). A classic example is hyperparameter optimization. Constrained optimization, on the other hand, involves solving the optimization problems under specific constraints (Bertsekas 2014; Goodfellow, Shlens, and Szegedy 2014). Examples include enforcing constraints of fairness, robustness, safety, etc. (Gallego-Posada 2024; Dwork et al. 2012). These optimization problems can be notoriously difficult to solve. Challenges arise from the complexity of formalizing constraints, the difficulty of creating differentiable relaxations, the absence of closed-form solutions, or evolving client requirements, among others.
@@
Interestingly, multiplicity offers a practical workaround: if you can’t optimize, search for it! Brute-force strategies of searching through various potential solutions are well-established in both bi-level optimization and constrained optimization literature (Sinha, Malo, and Deb 2017; Zhang et al. 2024; Gallego-Posada 2024). Though more efficient reductions and mathematical alternatives are preferred when feasible, they are not always possible. In such cases, searching through potential solutions to find the best fit becomes a viable strategy, and multiplicity plays a pivotal role in this search (see Figure 4). Multiplicity has been used in the literature to find models with lower bias (Black et al. 2024; Simson, Pfisterer, and Kern 2024; Ganesh et al. 2024; Islam, Pan, and Foulds 2021), smaller model complexity (Semenova, Rudin, and Parr 2022; Semenova et al. 2024; Boner et al. 2024; Rudin et al. 2024; Doctor, Mao, and Mhaskar 2024; Coupkova and Boutin 2024), better explanations (Sun, Wang, and Rudin 2024; Veran, Portier, and Fouquet 2023), improved generalizability (Li, Barnard, and Deng 2024), and the ability to allow personalization (Ma et al. 2020). Beyond simply enumerating the Rashomon set to search for better models, several recent works have also shown how the visualization of the Rashomon set can empower developers to select models that meet their specific requirements (He and Shaposhnik 2023; Eerlings et al. 2024; Luyten et al. 2024).
@@
Figure 4: The role of multiplicity in brute-force search for bi-level and constrained optimization problems in machine learning.
5.2 Ensembles and More
@@
Selecting the “best” model may not always be recommended, particularly when no single interpretation of the data can be optimal. Instead, combining insights from multiple models is preferred. Techniques like prediction ensembling, or bagging, have long been a central recommendation for stability in machine learning (Breiman 1996; Dietterich 2000; Long et al. 2024). Many methods in the multiplicity literature have capitalized on combining models in the Rashomon set. While literature in this direction primarily focuses on aggregating model explanations to create more stable and reliable explanations (Jiang et al. 2024a; Sreedharan, Kambhampati et al. 2018; Baniecki, Parzych, and Biecek 2024; Donnelly et al. 2023; Fisher, Rudin, and Dominici 2019; Dong and Rudin 2020; Kobylińska et al. 2024; Zuin et al. 2021; Hamamoto and Egi 2021; Kowal 2022), other works have also shown the benefits of aggregating fairness scores (Coston, Rambachan, and Chouldechova 2021), individual probabilities (Roth, Tolbert, and Weinstein 2023), or regression analysis to discover causality (Ueki and Kawasaki 2013). These techniques demonstrate the value of leveraging multiplicity not just to select a single best model but instead to combine multiple learned interpretations.
5.3 Hacking Metrics with Multiplicity
@@
While multiplicity can discover better models, it also introduces risks, particularly the potential of exploiting these methods to circumvent regulatory requirements and interventions. This is more prevalent when broad principles, such as fairness, are reduced to specific benchmarks or metrics (Black, Gillis, and Hall 2024). By leveraging multiplicity, it becomes possible to ‘hack’ these metrics, producing models that meet the specified criteria without truly adhering to the underlying principles (Coker, Rudin, and King 2021; Cooper et al. 2021; Møllersen and Holsbø 2023). Several studies in the literature have shown that such a search can produce models capable of “regulatory-washing”, being able to manipulate explanations (Sharma et al. 2024; A¨ıvodji et al. 2021; Shahin Shamsabadi et al. 2022) and fairness scores (Forde et al. 2021; Black, Gillis, and Hall 2024; Ganesh et al. 2024). Such manipulation can also occur unintentionally as a result of overfitting to a given metric (Black, Gillis, and Hall 2024; Cooper et al. 2021; Ganesh 2024), underscoring the need for vigilance against the misuse of multiplicity and calling for a more robust operationalization of regulatory frameworks (Black, Gillis, and Hall 2024).
6 Multiplicity and Responsible AI
@@
We now place multiplicity in the broader landscape of responsible AI. In this section, we examine the two key concerns for individuals originating from multiplicity, i.e., arbitrariness in model selection and outcome homogenization. Again, our discussion here builds on what was termed as the individual-level effects of multiplicity by Black, Raghavan, and Barocas (2022), while also focusing on topics that have gained more interest recently, such as homogenization.
6.1 Arbitrariness as a Responsible AI Concern
@@
Arbitrary decisions in an automated system can be deeply concerning when they can have direct and lasting impacts on human lives (Black, Raghavan, and Barocas 2022; Gomez et al. 2024; Sokol et al. 2024; Watson-Daniels et al. 2024). Borrowing an analogy from Gomez et al. (2024), imagine a judge deciding legal cases by flipping a coin. While this may seem extreme, it demonstrates how models can have arbitrariness embedded in them due to an analogous coin flip done by the developer during model design. This aligns closely with our ICA framework (§2.1), where we discuss the arbitrary choices that can contribute to multiplicity.
@@
There are contexts where a degree of “controlled randomness” may be acceptable—or even necessary (§6.2). However, arbitrariness is a significant concern in scenarios where individuals lack access to other ‘equivalent opportunities’ (Barocas, Hardt, and Narayanan 2023). For instance, in hiring, some level of arbitrariness may be acceptable, or even necessary to deal with the concerns of homogenization (§6.2). This is because individuals looking for a job often seek multiple opportunities across companies, increasing their chances of being hired elsewhere. In contrast, domains such as law or medicine typically involve singular, high-stakes decisions with no equivalent alternatives. In such situations, the presence of arbitrariness raises serious concerns for the responsible deployment of machine learning models.
@@
The negative effects of multiplicity extend beyond the arbitrariness of just the final prediction (§2.2). For instance, multiplicity in counterfactual explanations can impact the validity of algorithmic recourse (Hasan and Talbert 2022; Leventi-Peetz and Weber 2022; Baniecki, Parzych, and Biecek 2024; Hamman et al. 2023; Leofante, Botoeva, and Rajani 2023; Jiang et al. 2024a). The feasibility or nature of recourse might hinge on arbitrary design choices made during model development. These decisions can have real-world implications; for instance, a recourse provided by one model may become invalid if the model is updated, invalidating previous efforts. This inconsistency raises ethical and legal concerns (Kobylińska et al. 2024; De Toni et al. 2024).
Term Definition
Algorithmic Blackballing (Ajunwa 2020) ‘A worker’s lack of control over the portability of applicant data captured by automated hiring systems […] raising the specter of an algorithmically permanently excluded class of job applicants’
Algorithmic Monoculture (Kleinberg and Raghavan 2021) ‘The notion that choices and preferences will become homogeneous in the face of algorithmic curation’
Algorithmic Leviathan (Creel and Hellman 2022) ‘Automated decision-making systems that make uniform judgments across broad swathes of a sector.’
Outcome Homogenization (Bommasani et al. 2022) ‘The phenomenon of individuals (or groups) exclusively receiving negative outcomes from all ML models they interact with’
Generative Monoculture (Wu, Black, and Chandrasekaran 2024) ‘A distribution shift [towards less varied outputs] from source data (i.e., human-generated training data) to model-generated data (i.e., model outputs) for a specific task.’
Algorithmic Pluralism (Jain et al. 2024) ‘A state of affairs in which the algorithms used for decision-making are not so pervasive and/or strict as to constitute a severe bottleneck on opportunity.’
Table 2: Various terms used in the homogenization literature.
@@
This arbitrariness becomes even more problematic when it disproportionately impacts different individuals, particularly harming underrepresented demographics. As hinted earlier, a significant source of arbitrariness is the model’s lack of ability to learn the underlying distribution. Underrepresented groups often face these information gaps, which can manifest as data scarcity due to limited historical records or a lack of understanding of cultural context for how data relates to predictions (Barocas, Hardt, and Narayanan 2023; Hooker 2021; Buolamwini and Gebru 2018; Birhane 2022; Geirhos et al. 2020). Research has consistently shown that such disparities exacerbate existing inequalities—whether through arbitrariness, uncertainty, or multiplicity (Farnadi, Havaei, and Rostamzadeh 2024). The resulting disproportionate harms across groups highlight the pressing need to address arbitrariness in critical domains (Gomez et al. 2024; Cooper et al. 2024; Ali, Lahoti, and Gummadi 2021).
@@
Responsible AI Constraints and Multiplicity: We saw that multiplicity is a critical consideration in responsible AI. Interestingly, several works have also explored how multiplicity interacts with other pillars of responsible AI. Studies have shown that imposing fairness constraints can inadvertently increase multiplicity (Long et al. 2024; Cavus and Biecek 2024a). However, Long et al. (2024) argue that multiplicity stands outside the fairness-utility trade-off, meaning improvements in fairness do not have to entail increased multiplicity. They demonstrate that multiplicity can often be reduced through techniques like ensembling, while maintaining fairness. Similarly, Kulynych et al. (2023) explore the interaction between differential privacy and multiplicity, revealing that introducing privacy constraints tends to increase multiplicity. These findings underscore the complex interplay between multiplicity and other responsible AI principles, highlighting the need for further research to understand and navigate these trade-offs effectively.
6.2 Multiplicity and Homogenization